# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ 5.2. Software acceleration
Motivation. Deploying event cameras on mobile platforms requires software acceleration to handle sparse, asynchronous data in real time under strict resource limits. By transforming raw events into structured representations through clustering, compression, or sparse matrix operations, software acceleration streamlines denoising, filtering, feature extraction, and inference. Exploiting sparsity, reducing redundancy, and applying techniques like adaptive sampling and dynamic memory management are key to achieving energy-efficient, real-world mobile applications.
@@
Challenge. Effective software acceleration for event cameras on mobile platforms faces several challenges. First, sparse and asynchronous nature of event data makes dense-input algorithms inefficient, requiring specialized designs. Second, integration with deep learning frameworks is nontrivial, as most models are tailored for frame-based inputs. Third, limited on-device resources demand optimized memory management and throughput to balance speed and power. Finally, current benchmarks overlook unique characteristics of event data, calling for dedicated evaluation metrics.
@@
Literature review. In event-based vision systems, software acceleration is essential for improving the efficiency and performance of the event processing pipeline. By optimizing key stages of event handling, researchers can significantly enhance computational speed, reduce latency, and better utilize limited resources. Software acceleration techniques target various pipeline stages, such as event sampling, preprocessing, feature extraction, and event analysis. Each stage presents distinct challenges, and numerous specialized acceleration methods have been developed to address them:
@@
Event sampling. Event cameras capture raw data with precise timestamps and spatial locations. Unlike uniform sampling, adaptive sampling prioritizes significant scene changes to improve data efficiency. The key challenge is balancing temporal resolution with sparsity to prevent information loss or redundancy. [72] addresses this by introducing a recurrent convolutional SNN-based adaptive sampling module that dynamically adjusts rates based on spatio-temporal event patterns, enhancing overall efficiency.
@@
Event pre-processing. Events need to be denoised, filtered, and formatted to improve subsequent processing. Common challenges include false positives and inefficient filtering. Enhanced filtering algorithms, such as adaptive median filters, can retain important features while reducing computational overhead. A lightweight, hardware-friendly neural network architecture, 2-D CNN, is introduced in [176] for DVS gesture recognition, using a customized median filter to enhance signal-to-noise ratio and reduce hardware complexity.
@@
Feature extraction. Useful features, including spatial, temporal, and frequency features, are extracted from preprocessed data. Traditional techniques may not effectively capture event data’s unique properties. Software acceleration, particularly through deep learning models tailored for sparse data, can optimize feature extraction while reducing processing time. The FARSE-CNN model, proposed in [187], integrates sparse convolutional and asynchronous LSTM modules for efficient event data processing. [188] introduces SWformer, an attention-free architecture leveraging sparse wavelet transforms to capture high-frequency patterns, resulting in improved energy efficiency and performance.
@@
Event analysis. This step involves pattern recognition, classification, or regression, commonly for tasks like object detection or tracking. The sparsity of event data can lead to high computational costs. Optimized CNNs and SNNs can enhance the efficiency and effectiveness of analysis. A fast linear solver for camera motion restoration, developed in [189], addresses geometric problems and adapts to sudden motion changes, providing a robust solution for event data.
@@
Efficient event data representation. Efficient event data representation methods convert sparse, asynchronous event streams into structured formats for effective processing with reduced overhead. Recent approaches include event stacking [190], Temporal Activity Focus (TAF) [191], and HyperHistogram (HH) [192]. TAF adaptively adjusts time window length and resolution based on spatial and polarity cues, enhancing flexibility. HH builds multiple histograms from event polarity and temporal statistics, integrating them into 3D tensors to preserve fine-grained details.
@@
Geometric models-based methods. Geometric models leverage spatial relationships within scenes to improve the performance of event-based visual algorithms. By integrating geometric constraints from event data, these models reduce error accumulation, enhancing both accuracy and computational efficiency. Recent progress has been notable in event-based visual odometry [164], motion estimation [189, 193, 194, 195, 196], and time-to-collision estimation [197].
@@
Probabilistic models-based methods. Probabilistic models offer a principled way to represent data and quantify uncertainty, making them well-suited for the noisy, asynchronous nature of event streams. By leveraging probabilistic inference and optimization, they improve both robustness and accuracy in event-based vision. Recent studies demonstrate their effectiveness in object classification [198], where class likelihoods are directly maximized from event data, and in optical flow estimation [199], where noise modeling and belief propagation enhance precision.
@@
Deep learning-based methods. Deep learning has greatly advanced event-based vision by extracting complex spatio-temporal patterns for accurate prediction and decision-making. Recent efforts focus on efficiency, such as lightweight networks and sparse Transformers [200] for object detection, and deep models for dynamic obstacle avoidance [201] and tracking [202]. Advanced methods like GCNs and cross-representation distillation [203] further improve scene understanding, while point cloud-based networks enhance pose relocalization [132], and unified implicit neural representations support rolling shutter image restoration [204].
@@
A promising direction lies in lightweight neural architectures tailored to the sparse and asynchronous nature of event streams. Rather than compressing conventional models, these approaches co-design network structures and learning principles with the sensing modality itself. IDNet [144] replaces expensive 4D correlation volumes with an iterative motion-compensation loop, where a lightweight ConvGRU progressively refines residual flow, achieving competitive accuracy with far fewer parameters and memory. FARSE-CNN [187] proposes a fully asynchronous recurrent sparse-CNN, incorporating spatio-temporal compression modules to learn hierarchical features directly from events, yielding state-of-the-art efficiency at low computational cost. Ultralight Polarity-Split SNN [205] leverages polarity-split encoding and a learnable spatio-temporal loss for event-stream super-resolution, providing low-latency inference with minimal model size and enabling on-sensor deployment.
@@
Asynchronous and sparse computation-based acceleration. Asynchronous and sparse computation aligns with event cameras’ data characteristics by updating only when changes occur, selectively processing active regions. This reduces computational cost and boosts efficiency. Typical approaches focus on pixel brightness changes or region-specific features. Recent advances include sparse convolutional networks for asynchronous streams [206], graph-based frameworks [207], and local shift operations for optimized event handling [208].
6. Application: Mobile platform-based task
@@
This section explores the various tasks associated with mobile platforms, emphasizing the use of event cameras and other sensing technologies to improve performance in dynamic environments. These tasks are categorized into intrinsic sensing, external sensing, and event-based SLAM, each offering distinct applications and advantages, as shown in Fig.9.
6.1. Intrinsic sensing
@@
Vision odometry. Event-based visual odometry exploits asynchronous event streams to estimate motion with higher precision and lower computational cost than frame-based methods [209]. Early work focused on direct motion extraction, such as angular velocity estimation via contrast maximization on event edges [210] and improved attitude estimation under high rotation using enhanced aggregation functions [211]. More advanced systems integrate complementary modalities: [212] incorporates depth for robust odometry in challenging conditions, while first event-based stereo visual-inertial system [213] fuses stereo events, standard frames, and IMUs through spatiotemporal correlation and motion compensation, significantly boosting accuracy in dynamic scenes. Although these methods suggest proof-of-concept readiness, their large-scale deployment in unstructured real-world environments remains limited. Advanced methods have taken a step forward by attempting GPS-denied navigation for unmanned systems, comparing real-time terrain fingerprints generated from event camera outputs with pre-stored fingerprints derived from satellite imagery [214].
@@
Optical flow. Event cameras are highly suitable for optical flow estimation, supporting real-time processing in HDR environments critical for motion tracking. Existing methods fall into two main categories: (i) Deep learning-based approaches, which leverage neural networks to infer flow from event streams—for example, a hierarchical SNN for global motion perception [215], though they often require event-to-frame preprocessing. (ii) Traditional computer vision approaches, which exploit the spatio-temporal structure of event data. These include normal flow estimation fused with IMU data for high-dynamic velocity estimation [216] and contrast maximization methods with tailored reward functions [143]. Event-based optical flow has shown promise in drone status estimation, and deep learning methods perform well in lab settings; however, their robustness to sensor noise and illumination variability in real-world deployments remains an open research problem.
@@
Figure 9. Event camera-based sensing tasks on mobile platforms
6.2. External sensing
@@
Mapping. Event cameras improve 3D map construction by capturing rapid environmental changes with high temporal resolution. Feature-based methods, such as extracting points or lines, remain common—for instance, [217] detects straight lines in railway settings and combines them with odometry for infrastructure mapping. However, these methods depend on stable, repeatable features, which may be unreliable in sparse or occluded scenes. To address this, recent works adopt alternative strategies: BeNeRF [218] reconstructs NeRFs from a single blurred image and an event stream, reducing reliance on extensive pose data, while AsynHDR [219] leverages LCD modulation for HDR imaging, enriching scene information for more accurate 3D reconstruction. In summary, feature-based and neural approaches show promise for 3D reconstruction in dynamic environments, but many rely on assumptions about feature stability or controlled settings, indicating they are largely at the proof-of-concept stage rather than fully deployment-ready.
@@
Object detection & tracking. The high temporal resolution and low latency of event cameras make them particularly well-suited for precise object detection and tracking in dynamic environments [32, 220]. Existing methods can be broadly categorized into two groups: (i) Event stream-based detection and tracking methods. These approaches primarily leverage the spatio-temporal characteristics of event data. For example, [221] proposes a stereo event-based tracking algorithm that addresses occlusion by combining 3D reconstruction with cluster tracking. [222] proposes an end-to-end event cloud-based object tracking framework using density-insensitive key-event sampling, graph-based embedding, and motion-aware likelihood prediction. EVPropNet [223] tracks drone propellers using event data, while EDOPT [224] performs six-degree-of-freedom object pose tracking solely with event cameras. (ii) Fusion-based detection and tracking methods. These methods combine event data with additional sensors to enhance accuracy and robustness. [225] proposes a two-stage gaze estimation framework using event and frame data with anchor-based state shifts and denoising distillation. [226] explores object tracking from RGB and event data by leveraging a pre-trained ViT with mask modeling and orthogonal high-rank loss to enhance inter-modal token interaction. [227] proposes a motion-adaptive event sampling and bidirectional-enhanced fusion framework to align event and image data for more accurate object tracking. High-frequency drone localization using mmWave radar and event streams is demonstrated in [5], and [228] integrates depth camera data with events for obstacle tracking. In summary, some event-only methods and fusion-based approaches demonstrate impressive performance in dynamic and high-speed conditions. Fusion-based systems that combine RGB, depth, or radar data appear closest to real-world deployment, especially in drone navigation and obstacle tracking. In the industry, Meituan has explored combining event cameras with mmWave radar for UAV localization [5]. Tobii, a global leader in eye-tracking, together with Meta, has investigated event-camera-based gaze tracking [229, 230]. Meanwhile, SAAZ Micro Inc. and Neurobus have applied event cameras to drone detection [231, 232].
@@
Optical flow. Event cameras excel in external optical flow estimation, enabling robust motion tracking of dynamic objects in high-speed scenarios [233, 234]. Unlike internal perception, which focuses on self-motion estimation, external optical flow targets the movement of surrounding objects, supporting applications such as object capture and robotic control. Existing approaches can be broadly categorized into multimodal fusion, contrast maximization, and event-specific dense flow estimation. For instance, RPEFlow [235] enhances accuracy by fusing events with RGB images and point clouds via cross-modal attention. Contrast maximization (CM)-based methods [236, 237] provide a principled formulation and achieve state-of-the-art results with innovations like multi-reference focus loss and time-aware flow modeling. Complementarily, TEGBP [199] introduces an efficient framework for deriving dense optical flow directly from sparse event data. These developments highlight the growing versatility of event cameras for external motion analysis across diverse environments. Despite these advances, most methods remain at the proof-of-concept stage, validated primarily in controlled laboratory settings. It is worth noting that this technology shows significant potential for applications such as fall detection, crowd detection and tracking, and traffic data acquisition [238].
@@
Classification. Event cameras’ ability to capture rapid changes in lighting and motion greatly improves image classification [239, 240]. Event-based classification methods leverage subtle motion and illumination variations with specialized models. For instance, [241] uses graph convolutional networks on event-derived graph structures for classification, while [242] applies deep learning to recognize individuals from gait patterns. Event cameras also aid microscopic object classification using SNNs [243], and [244] captures facial micro-expressions for emotion recognition, showcasing event data’s sensitivity to subtle cues. These methods have been validated on specific datasets. While promising, their generalization to large-scale, unconstrained environments remains a challenge. It is worth noting that such approaches hold significant potential for applications in medical diagnostics [245].
@@
3D Reconstruction. Event cameras enable precise 3D reconstruction in dynamic scenes by leveraging high temporal resolution and sparse data. Real-time monocular reconstruction is shown in [246], while [247] used stereo event cameras for semi-dense 3D reconstruction, balancing accuracy and efficiency. Combining structured light with event cameras allowed high-speed 3D scanning with less data redundancy [248]. A polarization-based method in [249] maintains precision even at low event rates. Specific applications include real-time 3D hand gesture estimation [250], 3D human pose and shape estimation [251], and non-rigid object reconstruction handling complex motions [252]. In summary, real-time monocular and semi-dense stereo reconstructions have been achieved in controlled scenarios. Non-rigid object reconstruction and human pose estimation highlight potential, but deployment in complex scenes is still limited.
@@
Segmentation. Event cameras improve image segmentation performance by enabling accurate and efficient segmentation in rapidly changing environments [253, 254]. Segmentation approaches based on event data can be divided into: (i) Motion-based segmentation methods leverage the motion information captured by event cameras to segment scenes. For example, [255] proposes a motion compensation-based iterative optimization algorithm that segments scenes into independently moving objects, effectively exploiting the event camera’s high sensitivity to motion changes for dynamic target segmentation. (ii) Deep learning-based semantic segmentation methods utilize deep neural networks for feature extraction and semantic understanding. Existing works in this category explore how to effectively combine event data with RGB frames or adapt pre-trained segmentation models to the event domain. For instance, CMDA [256] exploits HDR of event cameras to complement the limited dynamic range of frame cameras, achieving robust semantic segmentation in challenging nighttime scenes. More recently, SAM-Event-Adapter [257] introduces a lightweight adapter to bridge event data with the Segment Anything Model (SAM), enabling zero-shot semantic segmentation and demonstrating strong generalization across diverse event datasets. Similarly, [258] proposes a multi-scale feature distillation method to align embeddings from event data with RGB images, further facilitating the adaptation of SAM for robust and universal object segmentation in the event domain. Although mostly evaluated on datasets, these methods have strong potential in IoT applications, such as crowd detection and traffic data acquisition, as well as in medical contexts like high-speed particle tracking in microfluidic devices, where accurate object segmentation is essential.

### v4

@@ 5.2. Software acceleration
Motivation. Deploying event cameras on mobile agents requires software acceleration to handle sparse, asynchronous data in real time under strict resource limits. By transforming raw events into structured representations through clustering, compression, or sparse matrix operations, software acceleration streamlines denoising, filtering, feature extraction, and inference. Exploiting sparsity, reducing redundancy, and applying techniques like adaptive sampling and dynamic memory management are key to achieving energy-efficient, real-world mobile applications.
@@
Challenge. Effective software acceleration for event cameras on mobile agents faces several challenges. First, sparse and asynchronous nature of event data makes dense-input algorithms inefficient, requiring specialized designs. Second, integration with deep learning frameworks is nontrivial, as most models are tailored for frame-based inputs. Third, limited on-device resources demand optimized memory management and throughput to balance speed and power. Finally, current benchmarks overlook unique characteristics of event data, calling for dedicated evaluation metrics.
@@
Literature review. In event-based vision systems, software acceleration is essential for improving the efficiency and performance of the event processing pipeline. By optimizing key stages of event handling, researchers can significantly enhance computational speed, reduce latency, and better utilize limited resources. Software acceleration techniques target various pipeline stages, such as event sampling, preprocessing, feature extraction, and event analysis, as shown in Fig.8. Each stage presents distinct challenges, and numerous specialized acceleration methods have been developed to address them:
@@
Event sampling. Event cameras capture raw data with precise timestamps and spatial locations. Unlike uniform sampling, adaptive sampling prioritizes significant scene changes to improve data efficiency. The key challenge is balancing temporal resolution with sparsity to prevent information loss or redundancy. [wang2025eas] addresses this by introducing a recurrent convolutional SNN-based adaptive sampling module that dynamically adjusts rates based on spatio-temporal event patterns, enhancing overall efficiency.
@@
Event pre-processing. Events need to be denoised, filtered, and formatted to improve subsequent processing. Common challenges include false positives and inefficient filtering. Enhanced filtering algorithms, such as adaptive median filters, can retain important features while reducing computational overhead. A lightweight, hardware-friendly neural network architecture, 2-D CNN, is introduced in [Fu2024A] for DVS gesture recognition, using a customized median filter to enhance signal-to-noise ratio and reduce hardware complexity.
@@
Feature extraction. Useful features, including spatial, temporal, and frequency features, are extracted from preprocessed data. Traditional techniques may not effectively capture event data’s unique properties. Software acceleration, particularly through deep learning models tailored for sparse data, can optimize feature extraction while reducing processing time. The FARSE-CNN model, proposed in [Santambrogio2024FARSE-CNN:], integrates sparse convolutional and asynchronous LSTM modules for efficient event data processing. [Fang2024Spiking] introduces SWformer, an attention-free architecture leveraging sparse wavelet transforms to capture high-frequency patterns, resulting in improved energy efficiency and performance.
@@
Event analysis. This step involves pattern recognition, classification, or regression, commonly for tasks like object detection or tracking. The sparsity of event data can lead to high computational costs. Optimized CNNs and SNNs can enhance the efficiency and effectiveness of analysis. A fast linear solver for camera motion restoration, developed in [Ren2024Motion], addresses geometric problems and adapts to sudden motion changes, providing a robust solution for event data.
@@
Efficient event data representation. Efficient event data representation methods convert sparse, asynchronous event streams into structured formats for effective processing with reduced overhead. Recent approaches include event stacking [8954323], Temporal Activity Focus (TAF) [10109007], and Hyper Histogram (HH) [Peng2023BetterAF]. TAF adaptively adjusts time window length and resolution based on spatial and polarity cues, enhancing flexibility. HH builds multiple histograms from event polarity and temporal statistics, integrating them into 3D tensors to preserve fine-grained details.
@@
Geometric models-based methods. Geometric models leverage spatial relationships within scenes to improve the performance of event-based visual algorithms. By integrating geometric constraints from event data, these models reduce error accumulation, enhancing both accuracy and computational efficiency. Recent progress has been notable in event-based visual odometry [Liu2021SpatiotemporalRF], motion estimation [Ren2024Motion, 10656961, wu2024motion, zhou2024resflow, zhou2025spatially], and time-to-collision estimation [li2024eventaidedtimetocollisionestimationautonomous].
@@
Probabilistic models-based methods. Probabilistic models offer a principled way to represent data and quantify uncertainty, making them well-suited for the noisy, asynchronous nature of event streams. By leveraging probabilistic inference and optimization, they improve both robustness and accuracy in event-based vision. Recent studies demonstrate their effectiveness in object classification [liu2020effectiveaerobjectclassification], where class likelihoods are directly maximized from event data, and in optical flow estimation [10204349], where noise modeling and belief propagation enhance precision.
@@
Deep learning-based methods. Deep learning has greatly advanced event-based vision by extracting complex spatio-temporal patterns for accurate prediction and decision-making. Recent efforts focus on efficiency, such as lightweight networks and sparse Transformers [peng2024sceneadaptivesparsetransformer] for object detection, and deep models for dynamic obstacle avoidance [sanket2020evdodgenetdeepdynamicobstacle] and tracking [10299598]. Advanced methods like GCNs and cross-representation distillation [deng2023dynamicgraphcnncrossrepresentation] further improve scene understanding, while point cloud-based networks enhance pose relocalization [ren2024simpleeffectivepointbasednetwork], and unified implicit neural representations support rolling shutter image restoration [lu2024uniinreventguidedunifiedrolling].
@@
A promising direction lies in lightweight neural architectures tailored to the sparse and asynchronous nature of event streams. Rather than compressing conventional models, these approaches co-design network structures and learning principles with the sensing modality itself. IDNet [wu2024lightweighteventbasedopticalflow] replaces expensive 4D correlation volumes with an iterative motion-compensation loop, where a lightweight ConvGRU progressively refines residual flow, achieving competitive accuracy with far fewer parameters and memory. FARSE-CNN [Santambrogio2024FARSE-CNN:] proposes a fully asynchronous recurrent sparse-CNN, incorporating spatio-temporal compression modules to learn hierarchical features directly from events, yielding state-of-the-art efficiency at low computational cost. Ultralight Polarity-Split SNN [xu2025ultralight] leverages polarity-split encoding and a learnable spatio-temporal loss for event-stream super-resolution, providing low-latency inference with minimal model size and enabling on-sensor deployment.
@@
Asynchronous and sparse computation-based acceleration. Asynchronous and sparse computation aligns with event cameras’ data characteristics by updating only when changes occur, selectively processing active regions. This reduces computational cost and boosts efficiency. Typical approaches focus on pixel brightness changes or region-specific features. Recent advances include sparse convolutional networks for asynchronous streams [messikommer2020eventbasedasynchronoussparseconvolutional], graph-based frameworks [9710541], and local shift operations for optimized event handling [10.1609/aaai.v37i2.25336].
6. Application: Mobile agent-based task
@@
This section explores the various tasks associated with mobile agents, emphasizing the use of event cameras and other perception technologies to improve performance in dynamic environments. These tasks are categorized into intrinsic perception, external perception, and event-based SLAM, each offering distinct applications and advantages (Fig.9).
6.1. Intrinsic perception tasks
@@
Vision odometry. Event-based visual odometry exploits asynchronous event streams to estimate motion with higher precision and lower computational cost than frame-based methods [10254473]. Early work focused on direct motion extraction, such as angular velocity estimation via contrast maximization on event edges [7805257] and improved attitude estimation under high rotation using enhanced aggregation functions [9706311]. More advanced systems integrate complementary modalities: [zuo2022devodeptheventcameravisual] incorporates depth for robust odometry in challenging conditions, while first event-based stereo visual-inertial system [10107754] fuses stereo events, standard frames, and IMUs through spatiotemporal correlation and motion compensation, significantly boosting accuracy in dynamic scenes. Although these methods suggest proof-of-concept readiness, their large-scale deployment in unstructured real-world environments remains limited. Advanced methods have taken a step forward by attempting GPS-denied navigation for unmanned systems, comparing real-time terrain fingerprints generated from event camera outputs with pre-stored fingerprints derived from satellite imagery [Neuromorphic].
@@
Optical flow. Event cameras are highly suitable for optical flow estimation, supporting real-time processing in HDR environments critical for motion tracking. Existing methods fall into two main categories: (i) Deep learning-based approaches, which leverage neural networks to infer flow from event streams—for example, a hierarchical SNN for global motion perception [8660483], though they often require event-to-frame preprocessing. (ii) Traditional computer vision approaches, which exploit the spatio-temporal structure of event data. These include normal flow estimation fused with IMU data for high-dynamic velocity estimation [lu2024eventbasedvisualinertialvelometer] and contrast maximization methods with tailored reward functions [Stoffregen2019EventCC]. Event-based optical flow has shown promise in drone status estimation, and deep learning methods perform well in lab settings; however, their robustness to sensor noise and illumination variability in real-world deployments remains an open research problem.
@@
Figure 9. Event camera-based perception tasks on mobile agents
6.2. External perception tasks
@@
Mapping. Event cameras improve 3D map construction by capturing rapid environmental changes with high temporal resolution. Feature-based methods, such as extracting points or lines, remain common. For instance, [9361194] detects straight lines in railway settings and combines them with odometry for infrastructure mapping. However, these methods depend on stable, repeatable features, which may be unreliable in sparse or occluded scenes. To address this, recent works adopt alternative strategies: BeNeRF [li2024benerfneuralradiancefields] reconstructs NeRFs from a single blurred image and an event stream, reducing reliance on extensive pose data, while AsynHDR [wu2024eventbasedasynchronoushdrimaging] leverages LCD modulation for HDR imaging, enriching scene information for more accurate 3D reconstruction. In summary, feature-based and neural approaches show promise for 3D reconstruction in dynamic environments, but many rely on assumptions about feature stability or controlled settings, indicating they are largely at the proof-of-concept stage rather than fully deployment-ready.
@@
Object detection & tracking. The high temporal resolution and low latency of event cameras make them particularly well-suited for precise object detection and tracking in dynamic environments [gehrig2024low, gehrig2023recurrentvisiontransformersobject]. Existing methods can be broadly categorized into two groups: (i) Event stream-based detection and tracking methods. These approaches primarily leverage the spatio-temporal characteristics of event data [ding2025hawkeye, chen2025count]. For example, [8088365] proposes a stereo event-based tracking algorithm that addresses occlusion by combining 3D reconstruction with cluster tracking. [zhu2022learning] proposes an end-to-end event cloud-based object tracking framework using density-insensitive key-event sampling, graph-based embedding, and motion-aware likelihood prediction. EVPropNet [sanket2021evpropnetdetectingdronesfinding] tracks drone propellers using event data, while EDOPT [10611511] performs six-degree-of-freedom object pose tracking solely with event cameras. (ii) Fusion-based detection and tracking methods. These methods combine event data with additional sensors to enhance accuracy and robustness. [zhu2025modeling] proposes a two-stage gaze estimation framework using event and frame data with anchor-based state shifts and denoising distillation. [zhu2023cross] explores object tracking from RGB and event data by leveraging a pre-trained ViT with mask modeling and orthogonal high-rank loss to enhance inter-modal token interaction. [chen2024crossei] proposes a motion-adaptive event sampling and bidirectional-enhanced fusion framework to align event and image data for more accurate object tracking. High-frequency drone localization using mmWave radar and event streams is demonstrated in [wang2025ultra], and [luo2024eventtracker] integrates depth camera data with events for obstacle tracking. In summary, some event-only methods and fusion-based approaches demonstrate impressive performance in dynamic and high-speed conditions. Fusion-based systems that combine RGB, depth, or radar data appear closest to real-world deployment, especially in drone navigation and obstacle tracking. In the industry, Meituan has explored combining event cameras with mmWave radar for drone localization [wang2025ultra]. Tobii, a global leader in eye-tracking, together with Meta, has investigated event-camera-based gaze tracking [tobii, stoffregen2022event]. Meanwhile, SAAZ Micro Inc. and Neurobus have applied event cameras to drone detection [saaz, neurobus].
@@
Optical flow. Event cameras excel in external optical flow estimation, enabling robust motion tracking of dynamic objects in high-speed scenarios [kosta2023adaptivespikeneteventbasedopticalflow, ponghiran2023eventbasedtemporallydenseoptical]. Unlike internal perception, which focuses on self-motion estimation, external optical flow targets the movement of surrounding objects, supporting applications such as object capture and robotic control. Existing approaches can be broadly categorized into multimodal fusion, contrast maximization, and event-specific dense flow estimation. For instance, RPEFlow [10376957] enhances accuracy by fusing events with RGB images and point clouds via cross-modal attention. Contrast maximization (CM)-based methods [Shiba_2022, 10517639] provide a principled formulation and achieve state-of-the-art results with innovations like multi-reference focus loss and time-aware flow modeling. Complementarily, TEGBP [10204349] introduces an efficient framework for deriving dense optical flow directly from sparse event data. These developments highlight the growing versatility of event cameras for external motion analysis across diverse environments. Despite these advances, most methods remain at the proof-of-concept stage, validated primarily in controlled laboratory settings. It is worth noting that this technology shows significant potential for applications such as fall detection, crowd detection and tracking, and traffic data acquisition [eventiot].
@@
Classification. Event cameras’ ability to capture rapid changes in lighting and motion greatly improves image classification [10377903, zheng2024eventdanceunsupervisedsourcefreecrossmodal]. Event-based classification methods leverage subtle motion and illumination variations with specialized models. For instance, [9337225] uses graph convolutional networks on event-derived graph structures for classification, while [9010397] applies deep learning to recognize individuals from gait patterns. Event cameras also aid microscopic object classification using SNNs [neftci2019surrogategradientlearningspiking], and [9844855] captures facial micro-expressions for emotion recognition, showcasing event data’s sensitivity to subtle cues. These methods have been validated on specific datasets. While promising, their generalization to large-scale, unconstrained environments remains a challenge. It is worth noting that such approaches hold significant potential for applications in medical diagnostics [medical].
@@
3D Reconstruction. Event cameras enable precise 3D reconstruction in dynamic scenes by leveraging high temporal resolution and sparse data. Real-time monocular reconstruction is shown in [Kim2016RealTime3R], while [Zhou_2018] used stereo event cameras for semi-dense 3D reconstruction, balancing accuracy and efficiency. Combining structured light with event cameras allowed high-speed 3D scanning with less data redundancy [Huang:21]. A polarization-based method in [muglikar2023event] maintains precision even at low event rates. Specific applications include real-time 3D hand gesture estimation [9711380], 3D human pose and shape estimation [zou2021eventhpeeventbased3dhuman], and non-rigid object reconstruction handling complex motions [xue2022eventbasednonrigidreconstructioncontours]. In summary, real-time monocular and semi-dense stereo reconstructions have been achieved in controlled scenarios. Non-rigid object reconstruction and human pose estimation highlight potential, but deployment in complex scenes is still limited.
@@
Segmentation. Event cameras improve image segmentation performance by enabling accurate and efficient segmentation in rapidly changing environments [jing2024hplesshybridpseudolabelingunsupervised, kong2024openesseventbasedsemanticscene]. Segmentation approaches based on event data can be divided into: (i) Motion-based segmentation methods leverage the motion information captured by event cameras to segment scenes. For example, [9010722] proposes a motion compensation-based iterative optimization algorithm that segments scenes into independently moving objects, effectively exploiting the event camera’s high sensitivity to motion changes for dynamic target segmentation. (ii) Deep learning-based semantic segmentation methods utilize deep neural networks for feature extraction and semantic understanding. Existing works in this category explore how to effectively combine event data with RGB frames or adapt pre-trained segmentation models to the event domain. For instance, CMDA [xia2023cmdacrossmodalitydomainadaptation] exploits HDR of event cameras to complement the limited dynamic range of frame cameras, achieving robust semantic segmentation in challenging nighttime scenes. More recently, SAM-Event-Adapter [yao2024sam] introduces a lightweight adapter to bridge event data with the Segment Anything Model (SAM), enabling zero-shot semantic segmentation and demonstrating strong generalization across diverse event datasets. Similarly, [chen2024segment] proposes a multi-scale feature distillation method to align embeddings from event data with RGB images, further facilitating the adaptation of SAM for robust and universal object segmentation in the event domain. Although mostly evaluated on datasets, these methods have strong potential in IoT applications, such as crowd detection and traffic data acquisition, as well as in medical contexts like high-speed particle tracking in microfluidic devices, where accurate object segmentation is essential.